I am having difficulty creating my own algebraic equation for class. I have to write an equation which models how old in years I will be when I am 3 times as old as my younger cousin. I am 33 she is 14.
here is what I have tried:
33x=19(3)
also I have tried
19x=33(3)
I am so confused right now.
I think I am going about this all wrong, can you just guide me in the right direction please. Thank you.

Found 2 solutions by ankor@dixie-net.com, checkley77:
Answer by ankor@dixie-net.com(22740)

write an equation which models how old in years I will be when I am 3 times as old as my younger cousin.
I am 33 she is 14.
:
Let y = no. of years for this to happen
:
33 + y = 3(14 + y)
33 + y = 42 + 3y
y - 3y = 42 - 33
-2y = 9
2y = -9
y = -4.5 yrs
:
this means you were 3 times as old as your cousin, 4.5 years ago
Will not happen in the future.
:
To answer the question; your age 4.5 yrs ago = 28.5; cousin was 9.5
:
Check:
33 - 4.5 = 3(14 - 4.5)
28.5 = 3(9.5)

Answer by checkley77(12844)

The formula is:
(33+x)=3(14+x)
33+x=42+3x
-2x=42-33
-2x=9
x=9/-2
x=-4.5 which says 4.5 years ago you were 3 times as old as your cousin.
Proof:
33-4.5=3(14-4.5)
28.5=3*9.5
28.5=28.5